OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (–8, 3); Slope: 6 (Points : 1) y + 3 = 6x – 48 y + 3 = 6(x + 8) y – 3 = 6(x – 8) y – 3 = 6(x + 8)

OpenStudy (anonymous):

The point slope formula is: y - y1 = m(x - x1). m = 6, y1 = 3, and x1 = -8 All we need to do is plug them in. y - 3 = 6(x + 8) <---- Answer. Since x1 is -8, -(-8) = 8. So we put an addition sign in front instead of the usual subtraction sign.

OpenStudy (anonymous):

thanks